The reaction A+B=C+D is studied in a one litre vessel at 250 degree celcius. The initial concentration of A was 3n and that of B was n. when eqilibrium was attained, equilibrium concentration of C was found to the equal to the equilibrium concentration of B. What is the concentration of D at equilibrium ?
 

Grade:11

2 Answers

Riddhi Jain
7 years ago
A + B = C + D3n n 0 0 at t=03n-x n-x x x at t=t`Since moles of B is equal to C at equilibriumTherefore, n-x=x So,x=n/2 A + B = C + D 3n-n/2 n-n/2 n/2 n/2Hence, D= n/2
ankit singh
3 years ago
The concentration of d at equilibrium will be n / 2.
Option (A) is correct.
Explanation:
The initial concentrations of the reactants are given below.
                             A   +     B   ⇄      C    +     D
At time "t" = 0    3n          n            0            0
At time "t" = 0    3n - x     n - x       x             x
n − x = x → x = n / 2
i.e. [D] equation  =n / 2
Thus the concentration of d at equilibrium will be n / 2.
Also learn more
What is the equilibrium concentration of each of the substances in the equilibrium when the initial concentration of ICl was 0.78 M?
2 ICl(g)⇌I2(g) + Cl2(g) ; KC = 0.14